Expand Fourier series f(x)=cos x limit -pi/2 to pi/2


1
Expert's answer
2021-06-02T12:33:43-0400
"f(x)=\dfrac{a_0}{2}+\displaystyle\sum_{n=1}^{\infin}\bigg(a_n\cos (\dfrac{n\pi x}{L})+b_n\sin (\dfrac{n\pi x}{L})\bigg)"

"a_0=\dfrac{1}{L}\displaystyle\int_{-L}^{L}f(x)dx"

"a_n=\dfrac{1}{L}\displaystyle\int_{-L}^{L}f(x)\cos (\dfrac{n\pi x}{L})dx"

"b_n=\dfrac{1}{L}\displaystyle\int_{-L}^{L}f(x)\sin(\dfrac{n\pi x}{L})dx"

"a_0=\dfrac{2}{\pi}\displaystyle\int_{-\pi/2}^{\pi/2}\cos x dx=\dfrac{2}{\pi}\big[\sin x\big]\begin{matrix}
 \pi/2 \\
 - \pi/2
\end{matrix}=\dfrac{4}{\pi}"

"a_n=\dfrac{2}{\pi}\displaystyle\int_{-\pi/2}^{\pi/2}\cos x\cos (2nx)dx"

"=\dfrac{2}{\pi}\big[-\dfrac{\cos((2n+1)x)}{2(2n+1)}-\dfrac{\cos((2n-1)x)}{2(2n-1)}\big]\begin{matrix}
 \pi/2 \\
 - \pi/2
\end{matrix}=0"

"=\dfrac{2}{\pi}\big[\dfrac{(-1)^n}{2n+1}-\dfrac{(-1)^n}{2n-1}\big]=\dfrac{4(-1)^{n+1}}{\pi(4n^2-1)}"


"b_n=\dfrac{2}{\pi}\displaystyle\int_{-\pi/2}^{\pi/2}\sin x\cos (2nx)dx=0"

"f(x)=\dfrac{2}{\pi}+\displaystyle\sum_{n=1}^{\infin}\big(\dfrac{4(-1)^{n+1}}{\pi(4n^2-1)}\cos (2nx)\big)"